# Mitigating Multimodal Hallucinations via Gradient-based Self-Reflection

### v2

@@ Abstract
Multimodal large language models achieve strong performance across diverse tasks but remain prone to hallucinations, where outputs are not grounded in visual inputs. This issue can be attributed to two main biases: text–visual bias, the overreliance on prompts and prior outputs, and co-occurrence bias, spurious correlations between frequently paired objects. We propose Gradient-based Influence-Aware Constrained Decoding (GACD), an inference-based method, that addresses both biases without auxiliary models, and is readily applicable to existing models without finetuning. The core of our approach is bias estimation, which uses first-order Taylor gradients to understand the contribution of individual tokens—visual features and text tokens—to the current output. Based on this analysis, GACD mitigates hallucinations through two components: (1) suppressing spurious visual features correlated with the output objects, and (2) rebalancing cross-modal contributions by strengthening visual features relative to text. Experiments across multiple benchmarks demonstrate that GACD effectively reduces hallucinations and improves the visual grounding of MLLM outputs.
@@ 1 Introduction
Recent advances in Multimodal Large Language Models (MLLMs) have demonstrated remarkable capabilities in producing coherent and context-aware content across a wide range of domains (Bai et al., 2023; Dai et al., 2023; Chen et al., 2024b; Liu et al., 2024a; Ye et al., 2024). Despite their impressive advancements, these models remain prone to hallucination, wherein the generated text is not faithfully grounded in the visual modality (Rohrbach et al., 2018; Li et al., 2023b). This limitation poses a critical barrier to establishing trust in the outputs of MLLMs.
@@
The hallucinations observed in MLLMs can be largely attributed to two fundamental biases (Kang & Choi, 2023; Li et al., 2023b; Kim et al., 2024). Text-visual bias refers to the excessive reliance on textual information—such as the input prompt and previously generated outputs—while neglecting the visual modality during generation. This bias becomes particularly pronounced in longer sequences, where MLLMs tend to depend more heavily on prior text and increasingly disregard visual cues (Zhou et al., 2023b; Favero et al., 2024). Co-occurrence bias arises from spurious statistical correlations embedded in the training data, which lead models to erroneously predict the presence of non-existent objects based on their frequent co-occurrence with observed objects in the visual inputs (Li et al., 2023b). This bias is particularly challenging to mitigate, and existing approaches largely rely on statistical priors rather than offering statistically agnostic solutions (Kang & Choi, 2023; Zhou et al., 2023a).
@@
Existing efforts to mitigate hallucinations in MLLMs can be broadly categorized into inference-based methods, which operate at the decoding stage (Chen et al., 2024a; Favero et al., 2024; Leng et al., 2024; Park et al., 2024; Woo et al., 2024), and training-based approaches, which intervene during model optimization (Ben-Kish et al., 2023; Chen et al., 2023b; Kang & Choi, 2023; Sun et al., 2023; Jiang et al., 2024; Yue et al., 2024). Inference-based approaches are valued for their cost-effectiveness, as they avoid the need for additional data collection, data bias examination, or extensive model retraining. However, these methods offer limited insight into the severity of underlying biases, leaving the root causes of hallucination insufficiently understood. In addition, some inference-based methods rely on auxiliary models—such as segmentation networks  Chen et al. (2024a), detection systems  Kan et al. (2024) , or even additional MLLMs (Radford et al., 2021; Deng et al., 2024; Xing et al., 2024)—which undermine their purported cost-effectiveness.
@@
Another limitation of existing methods lies in their lack of granularity when adjusting the underlying biases in MLLMs. Most approaches rely on heuristically tuned priors, which vary across datasets and fail to generalize reliably Leng et al. (2024); Zhao et al. (2024). Moreover, they apply uniform weighting across all visual features, offering no mechanism to selectively adjust bias at the level of individual featuresZhang et al. (2024b); Manevich & Tsarfaty (2024). This coarse treatment limits their effectiveness in mitigating co-occurrence bias, which arises from spurious statistical correlations between objects that are often represented by distinct visual features.
@@
  • Extensive experiments demonstrate that GACD mitigates hallucinations and enhances accuracy without sacrificing information. GACD achieves up to $8\%$ increase in overall score on AMBER Wang et al. (2023), an $8\%$ F1 boost on POPE Li et al. (2023b), up to $45\%$ improvement in detailness and a $92\%$ accuracy gain on LLaVA-QA90 Liu et al. (2024b).
@@ 2 Related Work
Hallucination and Bias. Hallucinations in LLMs often arise from biases in the training data McKenna et al. (2023); Huang et al. (2025), while in MLLMs, studies Tonmoy et al. (2024); Li et al. (2023b); Fu et al. (2024) show that hallucinations are closely linked to biases like text-visual and co-occurrence biases. Additionally, biases related to output position, which increase the risk of hallucination as output length grows, have been examined in Favero et al. (2024); Zhou et al. (2023b). Existing methods Li et al. (2023b); Fu et al. (2024); Kim et al. (2024) typically report only overall statistics, lacking a mathematical, sample-wise bias measurement. This distinction is important, as biases can vary case by case. Our approach measures sample-dependent bias via token-level gradient sensitivities, revealing how pre-trained MLLM parameters embed these biases  Kim et al. (2019); Guo et al. (2024), and enabling self-reflective hallucination mitigation.
@@
Hallucination Mitigation. Training-related hallucination mitigation methods Chen et al. (2023a); Jiang et al. (2024); Yue et al. (2024); Peng et al. (2025); Zadeh et al. (2025) are expensive, requiring access to training data and specialized statistical analysis. Among them, LPOI Zadeh et al. (2025) also employs an object-aware framework, highlighting the effectiveness of modeling object-level information for mitigating hallucinations. Reinforcement‐learning approaches Xing et al. (2024); Deng et al. (2024); Zhai et al. (2024) rely on supplementary feedback, often from human annotators or auxiliary LLMs/MLLMs, and the latter may themselves hallucinate. By contrast, post-decoding techniques modify model logits at inference time without further training or external feedback, making them lightweight add-ons. In text‐only LLMs, such methods aim to align outputs with factual knowledge Chuang et al. (2023); Li et al. (2023a). In MLLMs, post-decoding strategies emphasize the role of visual inputs Leng et al. (2024); Zhao et al. (2024); Favero et al. (2024) and can be classified into image-level and token-level interventions. Image-level decoding methods Zhang et al. (2024b); Manevich & Tsarfaty (2024) treat all objects in the input image uniformly, limiting their effectiveness in addressing co-occurrence hallucinations. Existing token-level methods either rely on external segmentation Chen et al. (2024a) and detection models Kan et al. (2024) or lack awareness of object-related decoupling Woo et al. (2024). Moreover, these methods typically introduce an implicit trade-off between accuracy and informativeness, reducing hallucinations at the expense of omitting valid details. Attention-based methods Tang et al. (2025); Zhang et al. (2024a) require careful selection of specific layers and often introduce model-specific adjustments or heuristics. In contrast, our GACD directly estimates embedded bias and decouples object-aware visual tokens, enabling sample-specific hallucination mitigation without external data, models, or model-specific adjustments, while achieving a more favorable balance between accuracy and informativeness.
@@ A MLLM with parameters $\theta$ computes, at each decoding step $m$, a logit vector
$\mathbf{z}_{m}\;=\;\pi_{\theta}\!\left(\mathbf{t}^{v},\,\mathbf{t}^{p},\,\mathbf{y}_{<m}\right)\in\mathbb{R}^{|\mathcal{V}|},\qquad\mathbf{y}_{<m}=[y_{1},\dots,y_{m-1}]\text{ (empty when }m=1\text{)}.$ (1)
This induces a categorical next-token distribution via the softmax $\sigma:\mathbb{R}^{|\mathcal{V}|}\!\to\!\Delta^{|\mathcal{V}|-1}$:
$p_{\theta}\!\left(y_{m}\mid\mathbf{t}^{v},\mathbf{t}^{p},\mathbf{y}_{<m}\right)\;=\;\big[\sigma(\mathbf{z}_{m})\big]_{y_{m}},\qquad 1\leq m\leq M,$ (2)
@@
Given a dataset $\mathcal{D}$ of $(\mathbf{t}^{v},\mathbf{t}^{p},\mathbf{y})$, maximum-likelihood training (or fine-tuning) estimates $\theta^{\star}$ by maximizing the conditional log-likelihood. Pretrained MLLMs encode statistical regularities (including spurious correlations) from training data in $\theta^{\star}$; such behavior can be probed without changing $\theta^{\star}$ via parameter-dependent analyses (e.g., gradients/attributions or counterfactual decodings) Kim et al. (2019); Guo et al. (2024), enabling self-reflective bias interpretation.
@@ 3.2 Gradient-Based Token Influence Estimation
To capture these embeded biases, we measure how each input token perturbs the output logits. Let $\mathbf{z}^{\star}_{m}\in\mathbb{R}^{|\mathcal{V}|}$ denote the step-$m$ logits $\mathbf{z}^{\star}_{m}=\pi_{\theta^{\star}}(\mathbf{t}^{v},\mathbf{t}^{p},\mathbf{y}_{<m})$. Around a reference sample point $(\mathbf{t}^{v(0)},\mathbf{t}^{p(0)},\mathbf{y}_{<m}^{(0)})$, the first-order Taylor expansion Spivak (1980) of the logits $\mathbf{z}^{\star}_{m}$ is
@@
where $\big|_{\cdot}$ indicate evaluation at the reference sample point. Taylor expansion details are in supplementary Sec. A. Each $\mathbf{g}^{v}_{ms},\,\mathbf{g}^{p}_{mn},\,\mathbf{g}^{y}_{mi}$ indicate a small token perturbation in its corresponding embedding space to a perturbation of the predict logit vector in $\mathbb{R}^{|\mathcal{V}|}$. We score the importance of each input token by the Manhattan norm of its gradient:
$I^{v}_{ms}\;=\;\|\mathbf{g}^{v}_{ms}\|_{1},\qquad I^{p}_{mn}\;=\;\|\mathbf{g}^{p}_{mn}\|_{1},\qquad I^{y}_{mi}\;=\;\|\mathbf{g}^{y}_{mi}\|_{1},$ (6)
@@
$\texttt{I}^{v}_{m}\;=\;\sum_{s=1}^{S}I^{v}_{ms},\qquad\texttt{I}^{p}_{m}\;=\;\sum_{n=1}^{N}I^{p}_{mn},\qquad\texttt{I}^{y}_{m}\;=\;\sum_{i=1}^{m-1}I^{y}_{mi}.$ (7)
@@ 3.3 Influence-Aware Constrained Decoding
GACD builds on token influence estimation with two components: (i) Object-aware Visual Token Grouping and (ii) Anchor-specific Influence-weighted Decoding. At step $m$, the former partitions visual tokens into object-related $\mathbf{t}^{o}$ and unrelated $\mathbf{t}^{u}$ based on objects detected in $\mathbf{y}_{<m}$. The latter extends contrastive decoding Li et al. (2022) by forming Anchor-specific negative guidance logits from pre-mentioned objects and computing a decoding weight $\alpha_{m}$ from token-influence measurements.
@@
$\mathcal{M}_{ms}=\mathbf{1}\!\left[\sum_{i=1}^{m-1}\mathcal{M}_{is}>0\right],\quad\mathcal{M}_{is}=\mathbf{1}\!\left[\,y_{i}\in\text{Noun}\;\wedge\;s=\arg\max_{j}I^{v}_{ij}\,\right],$ (8)
where $\mathbf{1}[\cdot]$ is the indicator and $\wedge$ is logical AND.
@@
$\mathbf{t}^{o}\;=\;\mathbf{t}^{v}\odot\mathcal{M}_{m},\qquad\mathbf{t}^{u}\;=\;\mathbf{t}^{v}\odot(\mathbf{1}-\mathcal{M}_{m}).$ (9)
@@
$\texttt{I}^{o}_{m}\;=\;\sum_{s=1}^{S}\|\mathbf{g}^{v}_{ms}\|_{1}\,\mathcal{M}_{ms},\qquad\texttt{I}^{u}_{m}\;=\;\sum_{s=1}^{S}\|\mathbf{g}^{v}_{ms}\|_{1}\,(1-\mathcal{M}_{ms}).$ (10)
@@
with $\alpha_{m}\geq 0$. In the probability space, moving along $\mathbf{z}^{\star}_{m}-\mathbf{z}^{o}_{m}$ increases the KL divergence $D_{\mathrm{KL}}(\sigma(\mathbf{z}^{\star}_{m})\|\sigma(\mathbf{z}^{o}_{m}))$ (see Sec. B). The original logits distribution $\mathbf{z}^{\star}_{m}$ can be viewed as $\pi_{\theta^{*}}(\mathbf{t}^{u},\mathbf{t}^{o},\mathbf{t}^{p},\mathbf{y}_{<m})$, i.e., a joint distribution that additionally depends on $\mathbf{t}^{u}$ compare to $\mathbf{z}^{o}_{m}$. Increasing the KL divergence therefore emphasizes the contribution of tokens $\mathbf{t}^{u}$, which are unrelated to previous mentioned objects, thereby mitigating co-occurrence bias in noun prediction. For non-noun steps, $\mathbf{t}^{u}$ coincides with $\mathbf{t}^{v}$, meaning that all visual tokens are emphasized. This adjustment helps reduce text–visual hallucination.
When analyzing token influence of $\hat{\mathbf{z}}_{m}$ in Eq. 11, the chain rule shows that $\mathbf{t}^{u}$ occur only in the original logits $\mathbf{z}^{\star}_{m}$ and are amplified by $(1+\alpha_{m})$, whereas other inputs also contribute to $\mathbf{z}^{o}_{m}$ and therefore undergo smaller influence changes. Let $\tilde{\texttt{I}}^{o}_{m},\tilde{\texttt{I}}^{p}_{m},\tilde{\texttt{I}}^{y}_{m}$ denote group influences computed on the negative branch $\mathbf{z}^{o}_{m}$ (analogous to equation 7). We then choose $\alpha_{m}$ so that the influence of $\mathbf{t}^{u}$ matches the dominant text level, $\texttt{I}^{t}_{m}\coloneqq\max(\texttt{I}^{p}_{m},\texttt{I}^{y}_{m})$. Aligning $\mathbf{t}^{u}$ influence with the question prompt $\texttt{I}^{p}_{m}$ is crucial for visually grounded responses, while balancing with previous outputs $\texttt{I}^{y}_{m}$ prevents visual forgetting.
@@
Unlike existing decoding methods Zhou et al. (2023b); Favero et al. (2024); Leng et al. (2024), which rely on adaptive plausibility constraints (e.g., prediction confidence) and require experimental tuning to determine optimal thresholds, our approach explicitly enforces non-negativity on the influence of object-related visual and prompt tokens. This corresponds to the following upper-bound condition:
@@
Additionally, since hallucinations are more likely in long generations Zhou et al. (2023b); Peng et al. (2025), we introduce a sample-dependent stopping criterion based on visual influence. Specifically, if the visual influence ratio $r^{v}_{m}$ of the token following the end-of-sequence (EOS) falls below a threshold $\epsilon$,

### v3

@@ Abstract
Multimodal large language models (MLLMs) achieve strong performance across diverse tasks but remain prone to hallucinations, where outputs are not grounded in visual inputs. This issue can be attributed to two main biases: text–visual bias, the overreliance on prompts and prior outputs, and co-occurrence bias, spurious correlations between frequently paired objects. We propose Gradient-based Influence-Aware Constrained Decoding (GACD), an inference-based method, that addresses both biases without auxiliary models, and is readily applicable to existing models without finetuning. The core of our approach is bias estimation, which uses first-order Taylor gradients to understand the contribution of individual tokens—visual features and text tokens—to the current output. Based on this analysis, GACD mitigates hallucinations through two components: (1) suppressing spurious visual features correlated with the output objects, and (2) rebalancing cross-modal contributions by strengthening visual features relative to text. Experiments across multiple benchmarks demonstrate that GACD effectively reduces hallucinations and improves the visual grounding of MLLMs outputs.
@@ 1 Introduction
Recent advances in Multimodal Large Language Models (MLLMs) show strong ability to produce coherent and context-aware content across a wide range of domains [bai2023qwen, dai2023instructblipgeneralpurposevisionlanguagemodels, chen2024internvl, liu2024llava, ye2024mplug]. Despite their impressive advancements, these models remain prone to hallucination, wherein the generated text is not faithfully grounded in the visual modality [rohrbach2018object, li2023evaluating]. This limitation poses a critical barrier to establishing trust in the outputs of MLLMs.
@@
The hallucinations observed in MLLMs can be largely attributed to two fundamental biases [kang2023impact, li2023evaluating, kim2024discovering]. Text-visual bias refers to the excessive reliance on textual information—such as the input prompt and previously generated outputs—while neglecting the visual modality during generation. This bias becomes particularly pronounced in longer sequences, where MLLMs tend to depend more heavily on prior text and increasingly disregard visual cues [zhou2023analyzing, favero2024multi]. Co-occurrence bias arises from spurious statistical correlations embedded in the training data, which lead models to erroneously predict the presence of non-existent objects based on their frequent co-occurrence with observed objects in the visual inputs [li2023evaluating]. This bias is particularly challenging to mitigate, and existing approaches largely rely on statistical priors rather than offering statistically agnostic solutions [kang2023impact, zhou2023solution].
@@
Existing efforts to mitigate hallucinations in MLLMs can be broadly categorized into inference-based methods, which operate at the decoding stage [chen2024halc, favero2024multi, leng2024mitigating, park2024convis, woo2024don], and training-based approaches, which intervene during model optimization [ben2023mocha, chen2023mitigating, kang2023impact, sun2023aligning, jiang2024hallucination, yue2024less]. Inference-based approaches are valued for their cost-effectiveness, as they avoid the need for additional data collection, data bias examination, or extensive model retraining. However, these methods offer limited insight into the severity of underlying biases, leaving the root causes of hallucination insufficiently understood. In addition, some inference-based methods rely on auxiliary models—such as segmentation networks  [chen2024halc], detection systems  [kan2024catchcomplementaryadaptivetokenlevel] , or even additional MLLMs [radford2021learning, deng2024seeing, xing2024efuf]—which undermine their purported cost-effectiveness.
@@
Another limitation of existing methods lies in their lack of granularity when adjusting the underlying biases in MLLMs. Most approaches rely on heuristically tuned priors, which vary across datasets and fail to generalize reliably [leng2024mitigating, zhao2024mitigating]. Moreover, they apply uniform weighting across all visual features, offering no mechanism to selectively adjust bias at the level of individual features [zhang2024debiasing, manevich2024mitigatinghallucinationslargevisionlanguage]. This coarse treatment limits their effectiveness in mitigating co-occurrence bias, which arises from spurious statistical correlations between objects that are often represented by distinct visual features.
@@
  • Extensive experiments demonstrate that GACD mitigates hallucinations and enhances accuracy while maintaining a favorable balance between accuracy and informativeness. GACD achieves up to $8\%$ increase in overall score on AMBER [wang2023llm], an $8\%$ F1 boost on POPE [li2023evaluating], up to $45\%$ improvement in detailness and a $92\%$ accuracy gain on LLaVA-QA90 [liu2024visual].
@@ 2 Related Work
Hallucination and Bias. Hallucinations in LLMs often arise from biases in the training data [mckenna2023sources, huang2025survey], while in MLLMs, studies [tonmoy2024comprehensive, li2023evaluating, fu2024mmecomprehensiveevaluationbenchmark] show that hallucinations are closely linked to biases like text-visual and co-occurrence biases. Additionally, biases related to output position, which increase the risk of hallucination as output length grows, have been examined in [favero2024multi, zhou2023analyzing]. Existing methods [li2023evaluating, fu2024mmecomprehensiveevaluationbenchmark, kim2024discovering] typically report only overall statistics, lacking a mathematical, sample-wise bias measurement. This distinction is important, as biases can vary case by case. Our approach measures sample-dependent bias via token-level gradient sensitivities, revealing how pre-trained MLLM parameters embed these biases  [kim2019learning, guo2024bias], and enabling self-reflective hallucination mitigation.
@@
Hallucination Mitigation. Training-related hallucination mitigation methods [chen2023minigpt, jiang2024hallucination, yue2024less, peng2025mitigating, zadeh2025lpoi] are expensive, requiring access to training data and specialized statistical analysis. Among them, LPOI [zadeh2025lpoi] also employs an object-aware framework, highlighting the effectiveness of modeling object-level information for mitigating hallucinations. Reinforcement‐learning approaches [xing2024efuf, deng2024seeing, zhai2024hallecontrolcontrollingobjecthallucination] rely on supplementary feedback, often from human annotators or auxiliary LLMs/MLLMs, and the latter may themselves hallucinate. By contrast, post-decoding techniques modify model logits at inference time without further training or external feedback, making them lightweight add-ons. In text‐only LLMs, such methods aim to align outputs with factual knowledge [chuang2023dola, li2023inference]. In MLLMs, post-decoding strategies emphasize the role of visual inputs [leng2024mitigating, zhao2024mitigating, favero2024multi] and can be classified into image-level and token-level interventions. Image-level decoding methods [zhang2024debiasing, manevich2024mitigatinghallucinationslargevisionlanguage] treat all objects in the input image uniformly, limiting their effectiveness in addressing co-occurrence hallucinations. Existing token-level methods either rely on external segmentation [chen2024halc] and detection models [kan2024catchcomplementaryadaptivetokenlevel] or lack awareness of object-related decoupling [woo2024don]. Moreover, these methods typically introduce an implicit trade-off between accuracy and informativeness, reducing hallucinations at the expense of omitting valid details. Attention-based methods [tang2025intervening, zhang2024redundancy] require careful selection of specific layers and often introduce model-specific adjustments or heuristics. In contrast, our GACD directly estimates embedded bias and decouples object-aware visual tokens, enabling sample-specific hallucination mitigation without external data, models, or model-specific adjustments, while achieving a more favorable balance between accuracy and informativeness.
@@ A MLLM with parameters $\theta$ computes, at each decoding step $m$, a logit vector
$\mathbf{z}_{m}\;=\;\pi_{\theta}\!\left(\mathbf{t}^{v},\,\mathbf{t}^{p},\,\mathbf{y}_{<m}\right)\in\mathbb{R}^{|\mathcal{V}|},$ (1)
where $\mathbf{y}_{<m}=[y_{1},\dots,y_{m-1}]\text{ (empty when }m=1\text{)}$. This induces a categorical next-token distribution via the softmax $\sigma:\mathbb{R}^{|\mathcal{V}|}\!\to\!\Delta^{|\mathcal{V}|-1}$:
$p_{\theta}\!\left(y_{m}\mid\mathbf{t}^{v},\mathbf{t}^{p},\mathbf{y}_{<m}\right)\;=\;\big[\sigma(\mathbf{z}_{m})\big]_{y_{m}},\quad 1\leq m\leq M,$ (2)
@@
Given a dataset $\mathcal{D}$ of $(\mathbf{t}^{v},\mathbf{t}^{p},\mathbf{y})$, maximum-likelihood training (or fine-tuning) estimates $\theta^{\star}$ by maximizing the conditional log-likelihood. Pretrained MLLMs encode statistical regularities (including spurious correlations) from training data in $\theta^{\star}$; such behavior can be probed without changing $\theta^{\star}$ via parameter-dependent analyses (e.g., gradients/attributions or counterfactual decodings) [kim2019learning, guo2024bias], enabling self-reflective bias interpretation.
@@ 3.2 Gradient-Based Token Influence Estimation
To capture these embeded biases, we measure how each input token perturbs the output logits. Let $\mathbf{z}^{\star}_{m}\in\mathbb{R}^{|\mathcal{V}|}$ denote the step-$m$ logits $\mathbf{z}^{\star}_{m}=\pi_{\theta^{\star}}(\mathbf{t}^{v},\mathbf{t}^{p},\mathbf{y}_{<m})$. Around a reference sample point $(\mathbf{t}^{v(0)},\mathbf{t}^{p(0)},\mathbf{y}_{<m}^{(0)})$, the first-order Taylor expansion [spivak1980calculus] of the logits $\mathbf{z}^{\star}_{m}$ is
@@
$\begin{split}\mathbf{g}^{v}_{ms}&\;\coloneqq\;\left.\frac{\partial\,\mathbf{z}^{\star}_{m}}{\partial\,\mathbf{t}^{v}_{s}}\right|_{\mathbf{t}^{v}=\mathbf{t}^{v(0)}},\qquad\mathbf{g}^{p}_{mn}\;\coloneqq\;\left.\frac{\partial\,\mathbf{z}^{\star}_{m}}{\partial\,\mathbf{t}^{p}_{n}}\right|_{\mathbf{t}^{p}=\mathbf{t}^{p(0)}},\\ \mathbf{g}^{y}_{mi}&\;\coloneqq\;\frac{\partial\,\mathbf{z}^{\star}_{m}}{\partial\,\mathbf{y}_{i}}\Big|_{\mathbf{y}=\mathbf{y}_{<m}^{(0)}},\end{split}$ (5)
where $\big|_{\cdot}$ indicate evaluation at the reference sample point. Taylor expansion details are in supplementary Sec. 1. Each $\mathbf{g}^{v}_{ms},\,\mathbf{g}^{p}_{mn},\,\mathbf{g}^{y}_{mi}$ indicate a small token perturbation in its corresponding embedding space to a perturbation of the predict logit vector in $\mathbb{R}^{|\mathcal{V}|}$. We score the importance of each input token by the Manhattan norm of its gradient:
$I^{v}_{ms}\;=\;\|\mathbf{g}^{v}_{ms}\|_{1},\quad I^{p}_{mn}\;=\;\|\mathbf{g}^{p}_{mn}\|_{1},\quad I^{y}_{mi}\;=\;\|\mathbf{g}^{y}_{mi}\|_{1},$ (6)
@@
$\texttt{I}^{v}_{m}\;=\;\sum_{s=1}^{S}I^{v}_{ms},\quad\texttt{I}^{p}_{m}\;=\;\sum_{n=1}^{N}I^{p}_{mn},\quad\texttt{I}^{y}_{m}\;=\;\sum_{i=1}^{m-1}I^{y}_{mi}.$ (7)
@@ 3.3 Influence-Aware Constrained Decoding
GACD builds on token influence estimation with two components: (i) Object-aware Visual Token Grouping and (ii) Anchor-specific Influence-weighted Decoding. At step $m$, the former partitions visual tokens into object-related $\mathbf{t}^{o}$ and unrelated $\mathbf{t}^{u}$ based on objects detected in $\mathbf{y}_{<m}$. The latter extends contrastive decoding [li2022contrastive] by forming Anchor-specific negative guidance logits from pre-mentioned objects and computing a decoding weight $\alpha_{m}$ from token-influence measurements.
@@
$\mathcal{M}_{ms}=\mathbf{1}\!\left[\sum_{i=1}^{m-1}\mathcal{M}_{is}>0\right],$ (8)
where $\mathcal{M}_{is}=\mathbf{1}\!\left[\,y_{i}\in\text{Noun}\;\wedge\;s=\arg\max_{j}I^{v}_{ij}\,\right]$ and $\mathbf{1}[\cdot]$ is the indicator and $\wedge$ is logical AND.
@@
$\mathbf{t}^{o}\;=\;\mathbf{t}^{v}\odot\mathcal{M}_{m},\quad\mathbf{t}^{u}\;=\;\mathbf{t}^{v}\odot(\mathbf{1}-\mathcal{M}_{m}).$ (9)
@@
$\texttt{I}^{o}_{m}\;=\;\sum_{s=1}^{S}\|\mathbf{g}^{v}_{ms}\|_{1}\,\mathcal{M}_{ms},\;\texttt{I}^{u}_{m}\;=\;\sum_{s=1}^{S}\|\mathbf{g}^{v}_{ms}\|_{1}\,(1-\mathcal{M}_{ms}).$ (10)
@@
with $\alpha_{m}\geq 0$. In the probability space, moving along $\mathbf{z}^{\star}_{m}-\mathbf{z}^{o}_{m}$ increases the KL divergence $D_{\mathrm{KL}}(\sigma(\mathbf{z}^{\star}_{m})\|\sigma(\mathbf{z}^{o}_{m}))$ (see Sec. 2 in supplementary). The original logits distribution $\mathbf{z}^{\star}_{m}$ can be viewed as $\pi_{\theta^{*}}(\mathbf{t}^{u},\mathbf{t}^{o},\mathbf{t}^{p},\mathbf{y}_{<m})$, i.e., a joint distribution that additionally depends on $\mathbf{t}^{u}$ compare to $\mathbf{z}^{o}_{m}$. Increasing the KL divergence therefore emphasizes the contribution of tokens $\mathbf{t}^{u}$, which are unrelated to previous mentioned objects, thereby mitigating co-occurrence bias in noun prediction. For non-noun steps, $\mathbf{t}^{u}$ coincides with $\mathbf{t}^{v}$, meaning that all visual tokens are emphasized. This adjustment helps reduce text–visual hallucination.
When analyzing token influence of $\hat{\mathbf{z}}_{m}$ in Eq. 11, the chain rule shows that $\mathbf{t}^{u}$ occur only in the original logits $\mathbf{z}^{\star}_{m}$ and are amplified by $(1+\alpha_{m})$, whereas other inputs ($\mathbf{t}^{o}$, $\mathbf{t}^{p}$, $\mathbf{y}_{<m}$) also contribute to $\mathbf{z}^{o}_{m}$ and therefore undergo smaller influence changes. Let $\tilde{\texttt{I}}^{o}_{m},\tilde{\texttt{I}}^{p}_{m},\tilde{\texttt{I}}^{y}_{m}$ denote group-level influences computed on the negative branch $\mathbf{z}^{o}_{m}$ (analogous to (7)). We then choose $\alpha_{m}$ so that the influence of $\mathbf{t}^{u}$ matches the dominant text level, $\texttt{I}^{t}_{m}\coloneqq\max(\texttt{I}^{p}_{m},\texttt{I}^{y}_{m})$. Aligning $\mathbf{t}^{u}$ influence with the question prompt $\texttt{I}^{p}_{m}$ is crucial for visually grounded responses, while balancing with previous outputs $\texttt{I}^{y}_{m}$ prevents visual forgetting.
@@
Unlike existing decoding methods [zhou2023analyzing, favero2024multi, leng2024mitigating], which rely on adaptive plausibility constraints (e.g., prediction confidence) and require experimental tuning to determine optimal thresholds, our approach explicitly enforces non-negativity on the influence of object-related visual and prompt tokens. This corresponds to the following upper-bound condition:
@@
Additionally, since hallucinations are more likely in long generations [zhou2023analyzing, peng2025mitigating], we introduce a sample-dependent stopping criterion based on visual influence. Specifically, if the visual influence ratio $r^{v}_{m}$ of the token following the end-of-sequence (EOS) falls below a threshold $\epsilon$,